# Probing metric fluctuations with the spin of a particle: a quantum simulation with bimodal optical cavities

### v1

@@ 3.1 Minimal model
where we have neglected the Fourier modes $\alpha_{m}(k)$ and $\beta_{m}(k)$ for $m\neq 0$ since they decouple from the fermonic modes. Out of all the possible momenta, the dominant terms in this Hamiltonian are the ones for which the resonant condition $\sqrt{2}\approx 2\pi k\mu$ holds, which defines a circle in momentum space of radius $k_{R}=\frac{1}{\sqrt{2}\pi\mu}$. Dropping the operator labels, the Hamiltonian can be rewritten as
$H=\sqrt{2}\sigma^{x}+\sqrt{2}\left(\alpha^{\dagger}\alpha+\beta^{\dagger}\beta\right)+g_{s-b}\left[\left(\beta+\beta^{\dagger}\right)\sigma^{x}+\left(\alpha+\alpha^{\dagger}\right)\sigma^{y}\right],$ (35)
$g_{s-b}=-\frac{\sqrt{2G}}{\sqrt{\pi}\mu^{3/2}}.$ (36)
4 Dynamical evolution of spin
We will investigate the behaviour of the spin under the influence of Hamiltonian (35) when it is initially prepared in various directions $x$, $y$ or $z$ and the graviton modes are in their ground state. From the bosonic part of (35) we see that the bosonic ground state corresponds to the zero-population Fock state |$\ket{0}_{\alpha}\otimes\ket{0}_{\beta}$. The initial state is thus:
@@ 4 Dynamical evolution of spin
|$\ket{\psi(t)}=e^{-i{H}t}\ket{\psi_{0}},$ (38)
for Hamiltonian (35) will in general give oscillatory dynamics between the spin and the bosons. Let us consider the case where $g_{s-b}=0$. Since the bosonic modes start in the ground state and they are decoupled from the spin, there are no dynamics in the bosonic populations. If the spin is also prepared in an eigenstate of $\sigma^{x}$ then no evolution will take place, while any other spin state will cause it to rotate due to the gravitational background classical field.
Let us now consider the case where $g_{s-b}\neq 0$. Even if the spin is prepared in the $x$-direction it will evolve as it is coupled non-trivially to the $\beta$-modes. Note that the evolutions corresponding to any other initial spin state will be asymmetrical due to the $\sigma^{x}$ self-energy of the spin. Due to the coupling, the spin-$x$ population is exchanged with the $\beta$-mode population and its oscillation amplitude will decay over time.
@@
In the following we will numerically investigate the time evolution for various coupling strengths. Note that $g_{s-b}$ depends on $G$ and $\mu$. For simplicity we take $\mu=1$, while $G$ is varied such that the regimes $G\ll\mu$, $G\sim\mu$ and $G\gg\mu$ are all considered. In the following section we determine numerically the time-evolution of the spin populations for different coupling regimes and different initial spin states.
4.1 Preparation in the spin-$x$ direction
@@
Figure 2: Time evolution of the spin and bosonic populations for weak coupling $G=0.05$, where the spin–boson interaction is much weaker than the bosonic self-interaction. The spin is initially prepared along the $x$-axis. The spin-$x$ population (solid) exhibits coherent exchange with the bosonic modes (dashed/dotted), while spin-$y$ and spin-$z$ remain unpopulated.
(a) Spin-$x$ dynamics for increasing coupling strength $G$. At $G=\pi$, the coupling equals $\sqrt{2}$ (the bosonic self-interaction strength). Stronger coupling leads to increased complexity in the spin dynamics.
@@
(b) Spin-$x$ population (horizontal axis: time, vertical axis: coupling $G$ on log scale). The white vertical line marks $G=\pi$ (coupling equal to self-interaction) separating the coherent from the incoherent time evolutions. The dashed black curve corresponds to the first peak for $G=0.05$ shown in Fig. 3(a).
Figure 3: Spin-$x$ population dynamics for the system initialized with spin aligned along the $x$-axis ($\mu=1$, matrix size $N=14$, see Hamiltonian (35)).
@@
To study the interplay between spin and bosonic modes, we initialise the system with the spin pointing along the $x$-axis and track its evolution under Hamiltonian (35). In the weak coupling regime ($G\ll 1$), where the spin–boson coupling is much smaller than the bosonic self-interaction strength, the spin-$x$ population exhibits regular, quasi-coherent oscillations, as shown in Fig. 2. These oscillations reflect the reversible exchange of excitation between the spin and the bosonic modes, with spin-$y$ and spin-$z$ remaining unpopulated due to the initial spin alignment and the decoupled dynamics.
However, even in this weak-coupling regime, small high-frequency fluctuations are visible on top of the main oscillations, indicating the onset of more complex dynamics. As the coupling strength $G$ increases, these micro-oscillations become more pronounced and ultimately dominate the dynamics, disrupting the coherent oscillation pattern.
@@
Fig. 3(b) illustrates how the spin-$x$ population evolves as a function of both time and coupling $G$ (plotted on a logarithmic scale). For $G\leq 0.1$, where the self-interaction dominates, the spin-$x$ population exhibits near-perfect revivals, indicating weak backaction from the bosonic environment. At $G=\pi$, the spin–boson coupling matches the bosonic self-interaction ($g_{s\text{-}b}=\sqrt{2}$), marking a crossover to a more complex regime. Beyond this point, revivals disappear, and the spin dynamics become increasingly irregular. For strong coupling ($G\geq 10$), the spin rapidly decoheres and fails to return to its initial value.
@@
This transition is more clearly visualised in Fig. 3(a), which shows time slices of the spin-$x$ dynamics for several values of $G$. As $G$ increases, two effects become prominent: (1) the emergence of fast oscillations superimposed on the slower envelope and (2) a decay in the amplitude of the envelope itself, signalling decoherence due to entanglement between the spin and the bosonic gravitational modes. Even though this decay is most visible in the bottom row of Fig. 3(a), it also appears at longer times in the upper panels, for example at $G=0.46$.
@@
4.2 Preparation in the spin-$y$ or spin-$z$ direction
@@
Fig. 4(a) shows the full population dynamics for weak coupling $G=0.05$. Initially, the spin is fully aligned along the $y$-axis, with zero population in spin-$x$ and spin-$z$. As time evolves, rapid oscillations develop between the spin-$y$ and spin-$z$ populations, reflecting precession in the $y$$z$ plane. The envelope of this oscillation exhibits slower modulations, corresponding to the gradual exchange of population with the bosonic modes. Meanwhile, the spin-$x$ population grows from zero and tracks this envelope, acting as a mediator between the spin subspace and the bosonic environment. The bosonic modes themselves remain nearly identical throughout the evolution due to the symmetry of their coupling.
@@
To highlight how the spin coherence degrades with increasing coupling strength, Fig. 4(b) plots the time evolution of the spin-$x$ and spin-$y$ populations for various values of $G$. At small $G$, clear oscillatory patterns are observed, with near-revivals of the initial spin-$y$ population. As $G$ increases, these revivals fade and are replaced by irregular micro-oscillations and a decay of the overall population envelope, signalling decoherence induced by stronger spin–boson entanglement. The system transitions from regular to chaotic dynamics around $G=\pi$, where the spin–boson coupling becomes comparable to the bosonic self-interaction.
Note that the spin-$z$ population, although not shown in Fig. 4(b), closely follows the spin-$y$ dynamics with a relative phase shift of $\pi/2$, consistent with the rotation symmetry of the model in the $y$$z$ plane.
@@
(a) Population dynamics for $G=0.05$. The spin-$y$ and spin-$z$ populations exhibit rapid oscillations with slowly varying envelopes. The spin-$x$ grows from zero and mediates coupling to the bosonic modes, which remain nearly identical.
(b) Time evolution of the spin-$x$ and spin-$y$ populations for different coupling strengths $G$. As $G$ increases, coherence is lost and the system transitions to chaotic dynamics. At $G=\pi$, the spin–boson coupling equals the bosonic self-interaction strength.
Figure 4: Spin and bosonic population dynamics for a spin initially aligned along the $y$-axis ($\mu=1$, matrix size $N=14$; see Hamiltonian (35)). The spin-$z$ population is not shown but closely follows spin-$y$ with a $\pi/2$ phase shift.
5 Optical cavity quantum simulation
@@
In a typical cavity-QED setup, an individual atom is strongly coupled to the quantised electromagnetic field inside a high-finesse cavity, where coherent interactions dominate over dissipative processes. To encode the two bosonic degrees of freedom $\alpha$ and $\beta$ appearing in Eq. (35) we can employ a bimodal cavity that supports simultaneously two nearly degenerate modes, distinguished, for example, by their polarisation or spatial structure. The atom itself provides the effective spin-$1/2$ system, with two long-lived hyperfine or Zeeman sublevels representing the $\ket{\uparrow}$ and $\ket{\downarrow}$ spin states.
@@
A concrete realisation can be achieved by simultaneously driving the atom and the cavity modes with appropriately tuned lasers. By choosing the laser amplitudes and detunings carefully, one can selectively couple each cavity mode to a specific atomic transition. One cavity mode may be chosen to drive spin rotations around the $x$-axis, while the other mediates rotations around the $y$-axis. The strength of these interactions is directly related to the effective spin-boson coupling parameter $g_{s-b}$, which in this setup can be tuned through the atom-cavity coupling strength, the detuning, and the drive amplitudes. Notably, such selective coupling strategies have been successfully employed to realise single- and two-photon blockades [24], as well as atom-photon state mapping and deterministic single-photon generation [25].
The initialisation of the system to the state (37) can be performed by optical pumping the atom to a well-defined electronic state that encodes the spin and by cooling the cavity modes close to their vacuum state, which corresponds to the bosonic ground states |$\ket{0}_{\alpha}\ket{0}_{\beta}$ in our model. The subsequent dynamical evolution of the spin can then be monitored by detecting the polarisation or temporal correlations of photons leaking from the cavity, providing direct access to the spin observables $\sigma_{i}(t)$ as well as the populations of the bosonic modes. State-of-the-art cavity-QED experiments already allow for the precise preparation, control, and readout of such hybrid atom-cavity systems with high fidelity and single-quantum resolution [24, 25]. Therefore, our proposed minimal model can be directly implemented with existing optical cavity technology, paving the way for controlled laboratory simulations of spin dynamics in fluctuating spacetime backgrounds.
@@ Acknowledgments
M.S. research projects are supported by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009). M.S. research is supported by the Ministry of Education, Singapore under the Academic Research Fund Tier 1 (FY2022, A-8000988-00-00). P.S.-R. has been supported by a Young Scientist Training Program (YST) fellowship at the Asia Pacific Center for Theoretical Physics (APCTP) through the Science and Technology Promotion Fund and the Lottery Fund of the Korean Government. P.S.-R. has also been supported by the Korean local governments in Gyeongsangbuk-do Province and Pohang City. This work was partially supported by the EPSRC grant EP/Z533634/1.

### v2

@@ 3.2 Dynamical evolution of spin
[…]
for Hamiltonian (35) will in general give oscillatory dynamics between the spin and the bosons. Let us consider the case where $g_{\text{s-b}}=0$: Since the bosonic modes start in the ground state and they are decoupled from the spin, there are no dynamics in the bosonic populations. If the spin is also prepared in an eigenstate of $\sigma^{x}$ then no evolution will take place, while any other spin state will cause it to rotate due to the gravitational background classical field.
Let us now consider the case where $g_{\text{s-b}}\neq 0$. Even if the spin is prepared in the $x$-direction it will evolve as it is coupled non-trivially to the $\beta$-modes. Note that the evolutions corresponding to any other initial spin state will be asymmetrical due to the $\sigma^{x}$ self-energy of the spin. Due to the coupling, the spin-$x$ population is exchanged with the $\beta$-mode population and its oscillation amplitude will decay over time.
@@
In the following we will numerically investigate the time evolution for various coupling strengths. Note that $g_{\text{s-b}}$ depends on $G$ and $\mu$. For simplicity we take $\mu=1$, while $G$ is varied such that the regimes $G\ll\mu$, $G\sim\mu$ and $G\gg\mu$ are all considered. In the following section we determine numerically the time-evolution of the spin populations for different coupling regimes and different initial spin states. In the numerical simulation we have to introduce a cut-off for the size of the otherwise infinite dimensional bosonic matrices $\alpha$ and $\beta$. Our numerical analysis shows that increasing the matrix size from $N=14$ to $N=15$ only changes the bosonic populations by at most $0.5\%$. We therefore consider $N=14$ to be a good approximation for these bosonic matrices in our simulation.
3.2.1 Preparation in the spin-$x$ direction
@@
Figure 2: Time evolution of the spin and bosonic populations for weak coupling $G=0.05$ ($\mu=1$), where the spin–boson interaction is much weaker than the bosonic self-interaction. The spin is initially prepared along the $x$-axis. The spin-$x$ population (solid) exhibits coherent exchange with the bosonic modes (dashed/dotted), while spin-$y$ and spin-$z$ remain unpopulated.
@@
Figure 3: Spin-$x$ population dynamics as a function of the coupling $G$ for the system initialized with spin aligned along the $x$-axis ($\mu=1$; see Hamiltonian (35)). At $G=\pi$, the coupling equals $\sqrt{2}$ (the bosonic self-interaction strength). (a) Spin-$x$ population (horizontal axis: time, vertical axis: coupling $G$ on log scale). The white vertical line marks $G=\pi$, roughly separating the coherent from the incoherent time evolutions. The dashed black curve corresponds to the first peak for $G=0.05$ shown in Fig. 3(b). (b) Spin-$x$ dynamics for increasing coupling strength $G$ (spin-$y$ is shown but does not have dynamics). Stronger coupling leads to increased complexity in the spin dynamics.
@@
To study the interplay between spin and bosonic modes, we initialise the system with the spin pointing along the $x$-axis and track its evolution under Hamiltonian (35). In the weak coupling regime ($G\ll 1$), where the spin–boson coupling is much smaller than the bosonic self-interaction strength, the spin-$x$ population exhibits regular, quasi-coherent oscillations, as shown in Fig. 2. These oscillations reflect the reversible exchange of excitation between the spin and the bosonic modes, with spin-$y$ and spin-$z$ remaining unpopulated due to the initial spin alignment and the decoupled dynamics. The frequency of the oscillations is controlled by the coupling $g_{\text{s-b}}$.
@@
The small high-frequency fluctuations, visible on top of the main oscillations, are due to the self-interactions of the bosons and the spin, with coupling $\sqrt{2}$, in (35). As the coupling strength $G$ increases, these micro-oscillations become more pronounced and ultimately dominate the dynamics, disrupting the coherent oscillation pattern. Fig. 3(a) illustrates how the spin-$x$ population evolves as a function of both time and coupling $G$ (plotted on a logarithmic scale). For $G\leq 0.1$, where the self-interaction dominates, the spin-$x$ population exhibits near-perfect revivals, indicating weak backaction from the bosonic environment. At $G=\pi$ (indicated by the white horizontal line in the figure), the spin–boson coupling matches the bosonic self-interaction ($g_{\text{s-b}}=\sqrt{2}$), marking a crossover to a more complex regime. Beyond this point, revivals disappear, and the spin dynamics become increasingly irregular. For strong coupling ($G\geq 10$), the spin rapidly decoheres and fails to return to its initial value.
@@
This transition is more clearly visualised in Fig. 3(b), which shows time slices of the spin-$x$ dynamics for several values of $G$. As $G$ increases, two effects become prominent: (i) the emergence of fast oscillations superimposed on the slower envelope and (ii) a decay in the amplitude of the envelope itself, signalling decoherence due to entanglement between the spin and the bosonic gravitational modes.
@@
3.2.2 Preparation in the spin-$y$ or spin-$z$ direction
@@
Fig. 4(a) shows the full population dynamics for weak coupling $G=0.05$. Initially, the spin is fully aligned along the $y$-axis, with zero population in spin-$x$ and spin-$z$. As time evolves, rapid oscillations develop between the spin-$y$ and spin-$z$ populations, reflecting precession in the yyzz plane. The envelope of this oscillation exhibits slower modulations, corresponding to the gradual exchange of population with the bosonic modes. Meanwhile, the spin-$x$ population grows from zero and tracks this envelope, acting as a mediator between the spin subspace and the bosonic environment. The bosonic modes themselves remain nearly identical throughout the evolution due to the symmetry of their coupling.
@@
To highlight how the spin coherence degrades with increasing coupling strength, Fig. 4(b) plots the time evolution of the spin-$x$ and spin-$y$ populations for various values of $G$. At small $G$, clear oscillatory patterns are observed, with near-revivals of the initial spin-$y$ population. As $G$ increases, these revivals fade and are replaced by irregular micro-oscillations and a decay of the overall population envelope, signaling decoherence induced by stronger spin–boson entanglement. The system transitions from regular to chaotic dynamics around $G=\pi$, where the spin–boson coupling becomes comparable to the bosonic self-interaction. Note that the spin-$z$ population, although not shown in Fig. 4(b), closely follows the spin-$y$ dynamics with a relative phase shift of $\pi/2$, consistent with the rotation symmetry of the model in the yyzz plane.
@@
Figure 4: Spin and bosonic population dynamics for a spin initially aligned along the $y$-axis ($\mu=1$; see Hamiltonian (35)). The spin-$z$ population is not shown but closely follows spin-$y$ with a $\pi/2$ phase shift. (a) Population dynamics for $G=0.05$. The spin-$y$ and spin-$z$ populations exhibit rapid oscillations with slowly varying envelopes. The spin-$x$ grows from zero and mediates coupling to the bosonic modes, which remain nearly identical. (b) Time evolution of the spin-$x$ (yellow line) and spin-$y$ (green envelope) populations for different coupling strengths $G$. As $G$ increases, coherence is lost and the system transitions to chaotic dynamics. At $G=\pi$, the spin–boson coupling equals the bosonic self-interaction strength.
3.2.3 Bloch sphere analysis
To characterise the spin’s evolution under its coupling with the bosonic modes, we define the Bloch vector $\boldsymbol{B}(t)=\big(\langle\sigma^{x}(t)\rangle,\;\langle\sigma^{y}(t)\rangle,\;\langle\sigma^{z}(t)\rangle\big)$, where expectation values are taken over the full system state. Figures 2, 3 and 4 display both the trajectories of the Bloch vector and the populations of the bosonic modes for spins initially prepared along different axes.
When the spin is prepared in the xx‑direction, $B_{y}$ and $B_{z}$ remain effectively zero. Hence the spin‐$x$ population (i.e. the probability of measuring “spin up in the xx‑basis”) coincides with the length of the Bloch vector $|\boldsymbol{B}|=\sqrt{\langle\sigma^{x}\rangle^{2}+\langle\sigma^{y}\rangle^{2}+\langle\sigma^{z}\rangle^{2}}$. In contrast, if the spin is initially along $y$, all three components evolve nontrivially. In Figure 5 (especially panel 5(a)) we show the Bloch‐sphere trajectory: the spin vector’s direction undergoes oscillation and its magnitude decays over time. This decay in $|\boldsymbol{B}|$ signals the decoherence of the spin subsystem due to entanglement with the bosonic modes. Under strong coupling (e.g. $G=10$) the decay is rapid. After an initial drop, the magnitude tends to stabilise on average, with residual micro‐oscillations reflecting finite‐size effects of the bosonic sector.
To explore how decoherence scales with the coupling strength, we plot the time‐averaged spin‐vector magnitude in Figure 5(b) as a function of $G$. For small couplings, oscillations of the Bloch vector are regular and revivals of $|\boldsymbol{B}|$ are visible, although the average magnitude is reduced compared to the ideal isolated spin. As $G$ increases, the dynamics becomes more irregular (“chaotic” in appearance), revivals are suppressed, and the time‐averaged magnitude is much smaller. For very large coupling values we even observe a slight increase in the average amplitude compared to immediate large decays, which we attribute to the fact that only a finite number of bosonic modes are resonantly coupled — the system is not truly in the infinite‐bath limit, so complete decoherence (i.e. $|\boldsymbol{B}|\to 0$) does not strictly occur.
A useful comparative reference is the well-known “spin‐boson model”, which describes a two‐level system coupled to an environment of harmonic oscillators. In that model, dephasing and loss of coherence (seen as shrinking Bloch‐vector length) are standard features and have been studied both analytically and numerically. For example, Orth et al. [22] studied nonperturbative decoherence in driven spin‐boson systems, observing suppression of coherence and decay of Bloch‐vector magnitude as coupling increases.
Figure 5: Evolution of the Bloch vector $\boldsymbol{B}(t)=(\langle\sigma^{x}\rangle,\langle\sigma^{y}\rangle,\langle\sigma^{z}\rangle)$ for a spin initially prepared along the $y$-axis ($\mu=1$; dynamics governed by the Hamiltonian in eq. (35)). (a) The spin trajectory on the Bloch sphere for coupling $G=10$. The vector precesses primarily in the yyzz plane while its length gradually decreases, indicating decoherence due to entanglement with the bosonic modes. (b) Time-averaged Bloch vector length as a function of coupling strength $G$. For illustration, the vector length $|\boldsymbol{B}(t)|$ is also shown as a function of time, for two representative couplings, $G=0.05$ and $G=100$ (showing the time-averaged value as a horizontal line), highlighting the transition from weak to strong decoherence.
3.3 Optical cavity quantum simulation
@@
[…]
@@
A concrete realisation can be achieved by simultaneously driving the atom and the cavity modes with appropriately tuned lasers. By choosing the laser amplitudes and detunings carefully, one can selectively couple each cavity mode to a specific atomic transition. One cavity mode may be chosen to drive spin rotations around the $x$-axis, while the other mediates rotations around the $y$-axis. The strength of these interactions is directly related to the effective spin-boson coupling parameter $g_{\text{s-b}}$, which in this setup can be tuned through the atom-cavity coupling strength, the detuning, and the drive amplitudes. Notably, such selective coupling strategies have been successfully employed to realise single- and two-photon blockades [25], as well as atom-photon state mapping and deterministic single-photon generation [26].
[…]
@@ Acknowledgments
M.S. research projects are supported by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009). M.S. research is supported by the Ministry of Education, Singapore under the Academic Research Fund Tier 1 (FY2022, A-8000988-00-00). P.S.-R. has been supported by a Young Scientist Training Program (YST) fellowship at the Asia Pacific Center for Theoretical Physics (APCTP) through the Science and Technology Promotion Fund and the Lottery Fund of the Korean Government. P.S.-R. has also been supported by the Korean local governments in Gyeongsangbuk-do Province and Pohang City. This work was partially supported by the EPSRC grants EP/Z533634/1 and UKRI1337:Anyons24.